OpenStudy (anonymous):

1. 1/x divided by 7 2. y^2-49/y+7 divided by y-7/8 3. 6/2a-5 divided by 8/5-2a

OpenStudy (anonymous):

\[1)\space \frac{1}{7x}\] \[2)\space {8}\] \[3)\space \frac{-6}{8} \space=\frac{-3}{4}\]